--dont touch table-->
The Weeks of February 16-26, 2004
Last week, workmen continued insulating the sub-basement and third floor, and they sealed ducts and grouted dampers. Electricians pulled wires in the H & R building and RC1, and they installed temporary lights in the sub-basement. Plaster trim work on the ceiling continued and the areaway was insulated. Workmen continued working on the northwest flat roof area installing scratch coating. Trench work started in preparation for laying 8" pipes and the 8" water vault was also worked on. Steam-chill water pipes were roughed-in and wires were pulled in the sub-basement. The mop sink in the janitor's closet area was framed also. The flat roof work was completed earlier in the work. On Wednesday, steam/chill water vent line manhole # 1 was worked on and work in the sub-basement continued also. Work on the northwest tower scratch work continued. The valley plate and roof panels were installed on the south side of the building. Outside, workmen continued laying 8" and 4" pipes west to east on the Adams St. side of the building. Roof panels were installed and workmen also backfilled over the exterior water pipe work in back of the building. A pad was poured for manhole # 2. This week, work continued on both the sub-basement and third floor ducts. Electricians pulled wiring in the basement and hooked up wires. Work began for installing the starter in the H & R building. Metal flashing was also installed in the upper level troughs. Later, wiring was hooked up in the transformer vault and wires were also pulled to panels. Sheetrock was hung in the electrical room. Patch work started to repair leaks in the roof found after heavy rains. Backfill work also continued outside and a top vault was set in place. Trench work then started on the opposite side of the driveway. Sheet rock was hung in the fire corridor on the third floor along with the chase area.
The Weeks of February 2-13, 2004
On Monday, 2 February, contractors worked on the interior lobby columns, hatchannel in the research room, and trim in the high ceiling mezzanine. Electricians ran pipe in RC1 and pulled wire in the sub-basement. Scratch coat continued on the exterior north side of the building and casting crates for the building were unloaded. Metal flashing was also installed on the north side of the building. Mezzanine handrail work continued and the RC1 door was grouted. On Tuesday, workmen tied in the damper on the 2nd floor and 1st floor. Pull string was installed in pipe by electricians. The ceiling was framed in the research room and workmen worked on trim in the high ceiling. Scratch work continued on the building exterior. Ridge and valley tin was installed. Water was pumped out of the building after rains. On Wednesday, sub-basement and 1st floor controls were worked on and workmen insulated the sub-basement and 1st floor also. Pipe continued to be run in RC1. Plaster trim work began in the hi-mezzanine ceiling and sheetrock trim work was performed on the 1st floor. On the building exterior, cornices were installed and scratch coat was also installed. Valley and ridge tin work continued along with metal flashing. Tarps were installed on the flat roof and wood block work also continued on the roof. On Thursday, workmen worked on the copper controls in the basement along with insulation punch work on the 2nd floor, 1st floor and sub-basement. Wires continued to be pulled in RC1, and reels were unloaded. High ceiling plaster and sheetrock trim continued in the reference room, along with cornice work on the building exterior north side. Pads were poured in the sub-basement. On Friday, wires were pulled in RC1 and plaster trim work continued in the high ceiling area and copper continued in the basement. Floors were swept of water due to rains again. Outside the building, workmen put out several yards of gravel to halt erosion caused by the rains along with the construction work.
Updated: March 2, 2003
Alabama Department of Archives & History
624 Washington Avenue
Montgomery, Alabama 36130-0100
Phone: (334) 242-4435